Question: Your older cousin decides to gift you U.S. $100 in fiat currency in a year from now as a college graduation gift. She is open to the idea of giving you the gift today to help you pay some of your outstanding bills. How much money would you be willing to accept if it was offered today instead of one year from now?

Response, in $: M Your personal discount rate, r = (100 – M)/M You may directly enter any other feasible discount rate with appropriate justification.
Answer #1

Let personal discount rate be 1%, the return I can earn if I keep the money in bank

Present Value=Future Value/(1+r)^t=100/(1+1%)^1=99.0099010